A sample size 25 is picked up at random from a population which is normally distributed with a mean of 100 and variance of 36. Calculate:-

(a) Pr{ ͞X≤99}

(b) Pr{98≤ ͞X≤100}

1
Expert's answer
2021-03-23T11:21:08-0400

Using the formula;


a)

= 100+ 0.99(362/ 5) = 356.61

b) 100+ 0.98, 1.0 (36/5) = 364.3, 359.2